OpenStudy (anonymous):

Initially I had $45.00 but then I ended up with $37.50, now I'm completely confused. Please help. What is the simple interest to the nearest cent on a credit card balance of $1,250 at 12% interest for 3 months? A) $38.50 B) $45.00 C) $37.50 D) $46.00

OpenStudy (amistre64):

how did you work it to start with? the formula is pretty easy to follow

OpenStudy (amistre64):

interest tends to be stated for a unit of time as a year ... so 12 months. the time factor will be 3/12 which is about the only thing that might be tricky with it

OpenStudy (amistre64):

IF we had erred and said t=3 instead of t=3/12 we would have interest of $450 not $45

OpenStudy (amistre64):

12% in 12 months is 1% a month ... or 3% in 3 months 12.50 = 1% ... multiply by 3 for the dbl chk

OpenStudy (anonymous):

olay, I was doing 1250x (1+0.12)^3/12

OpenStudy (amistre64):

you compounded then .... which is related to simple interest

OpenStudy (anonymous):

I'm wondering if I did just 3 in error so I had the first answer.

OpenStudy (amistre64):

simple interest = Prt 1250 x .12 x 3/12

OpenStudy (amistre64):

if we add the interest back to the prinicipal .. A = P + Prt A = P(1 + rt) ; compounding adds interest to added interest ...

OpenStudy (anonymous):

So since I'm just finding the what the interest is. I just use Prt & the answer is 37.50?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

That makes sense now. thank you so much!

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

Have a great day! :)